# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

### v1

@@ Appendix B Technical Details for MaxCut Hamming and MaxBisection
$-\mathbb{E}_{\pi}H(x)=\sum_{i<j}e_{ij}\mathbb{E}_{\pi}\frac{1}{2}(1-x_{i}x_{j})=\sum_{i<j}e_{ij}\Pr[x_{i}\neq x_{j}].$
[…]
@@
$-\frac{\mathbb{E}_{\pi}H(x)}{n}=\frac{p}{4}(1+o(1))$ (49)
@@
Fix a graph $G$, and consider an arbitrary $x\in\{-1,1\}^{n}$ such that $\lvert x\rvert=k$ corresponding to number of $+1$’s. Swaps only occur between $x_{i}$ and $x_{j}$ that are different. All one step transitions denoted by $\sim$ are implied to be under the transposition walk.
@@
For a given fixed $x\in\{-1,1\}^{n}$ satisfying $\lvert x\rvert=k$, we know there are $k(n-k)$ pairs of indices such that $x_{i}\neq x_{j}$, $\binom{n-k}{2}$ indices where $x_{i}=x_{j}=-1$, and $\binom{k}{2}$ indices where $x_{i}=x_{j}=1$. Thus we can decompose the expectation as follows:
@@
[…]
@@
    ±$\pm 1:\mathbb{E}_{y\sim x}\left[\frac{1-y_{i}y_{j}}{2}\frac{1-y_{r}y_{s}}{2}\right]=\frac{4}{(n-k)k}$
@@
[…]
Then we compute $-H(x)[2\mathbb{E}_{y\sim x}[H]-H(x)]$:
@@ Appendix C Constrained Short Path via Penalized Objective
The Markov Chain $\mathcal{M}=(\mathcal{X},P,\pi)$ is the random walk on the hypercube, and hence $\pi$ is the uniform distribution over $\{-1,1\}^{n}$. Thus, the short-path Hamiltonian is the same as [DPCB23] but with the penalized cost Hamiltonian.
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be random walk on the $n$-bit Hamming cube. Let $H:\left\{-1,1\right\}^{n}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy
@@
If the number of constraints is $m=\Theta(n)$, then there exists a short-path algorithm with runtime
@@
The penalty method ensures that our algorithm only outputs feasible solutions, but the framework only provides a speedup over unconstrained bruteforce search. Hence, the generalized short-path algorithm is significantly more effective.
@@ D.2 Ground State Preparation for Glauber Mixer
|$\ket{\pi^{(2)}}=\sum_{x\in\mathcal{X}}\sqrt{\pi_{\lambda_{2}}(x)}\ket{x}.$
This quantum state can be prepared by applying ground state projector of $D_{\lambda_{1}}$ to and $D_{\lambda_{2}}$ to |$\ket{\pi^{(0)}}$ through fixed point amplitude amplification. We repeat this process until we prepare |$\ket{\pi^{(k)}}$,
|$\ket{\pi^{(k)}}=\sum_{x\in\mathcal{X}}\sqrt{\pi_{\lambda_{k}}(x)}\ket{x}$
@@
Let $\delta(\lambda)>0$ be the spectral gap of the discriminant matrix $D(P_{\lambda})$ associated with $P_{\lambda}$. Then, there exists a quantum algorithm that prepares the ground state of $-D$ up to $\varepsilon$ accuracy in $L2$ norm with run time $\widetilde{\mathcal{O}}(\delta_{\textup{min}}^{-1/2}\log(1/\varepsilon))$ where $\delta_{\textup{min}}=\inf_{0\leq\lambda^{\prime}\leq\lambda}\delta(\lambda^{\prime})$.
@@
|$\ket{\pi^{(0)}}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\ket{x_{i}}$
where $x_{i}$ is all 00 bit string except location $i$. This quantum state is ground state of $D$ at $\lambda=0$ as each $x_{i}$ is an independent set and they are the only ones with non-zero probability. This quantum state is essentially superposition over all strings with Hemming weight 1. This state can be prepared by applying $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{i}$ where $X_{i}$ is Pauli-$X$ applied to the all 00 state. Therefore, this state can be prepared in polynomial time. However, we cannot implement the Glauber dynamics at $\lambda=0$ as the transition density is not meaningful. Instead, we start with |$\ket{\pi^{(1)}}$ which is the ground state of $D_{\lambda_{1}}$ which can be prepared from |$\ket{\pi^{(0)}}$ since |$\ket{\pi^{(0)}}$ and |$\ket{\pi^{(1)}}$ overlaps significantly
$|\braket{\pi^{(0)}}{\pi^{(1)}}|=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{\lambda_{1}^{1/2}}{\sqrt{Z_{1}}}\geq\frac{\sqrt{n\lambda_{1}}}{\sqrt{n\lambda_{1}+\sum\limits_{|\mathcal{I}|>1}\lambda_{1}^{\mathcal{|I|}}}}$ $=\Omega(1),$
for $\lambda_{1}\leq\frac{3n}{2^{n}}$. Similarly, given |$\ket{\pi^{(k-1)}}$, we can prepare |$\ket{\pi^{(k)}}$ efficiently. The number of calls to the ground state projector by the fixed point amplitude amplification from |$\ket{\pi^{(k-1)}}$ to |$\ket{\pi^{(k)}}$ is $\widetilde{\mathcal{O}}(|\langle\pi^{(k-1)}|\pi^{(k)}\rangle|^{-1})$. The overlap can be calculated as
$\left|\!\braket{\pi^{(k-1)}}{\pi^{(k)}}\!\right|$ | $=\sum_{x\in\mathcal{X}}\frac{\lambda_{k-1}^{|x|/2}}{\sqrt{Z_{k-1}}}\frac{\lambda_{k}^{|x|/2}}{\sqrt{Z_{k}}}$
@@
if we set $\Delta\leq\frac{2}{n}$. Hence, starting from |$\ket{\pi^{(1)}}$ we can prepare a schedule that maintains constant overlap with the subsequent state. Since we have constant overlap throughout the schedule, each amplitude amplification step only requires constant number of calls to ground state projectors. Also note that implementing the ground state projector requires $\widetilde{\mathcal{O}}(\delta^{-1})$ calls to block-encoding of $D$ [GSLW19]. Finally, we only need to do $\mathcal{O}(\operatorname{poly}(n))$ rounds of amplitude amplification since $\lambda_{k}=\lambda_{1}\exp(2k/n)$ and for $k\geq\frac{n}{2}\log(\lambda_{c}/\lambda_{1})$, $\lambda_{k}\geq\lambda_{c}$.
@@
Let $\delta(\beta)>0$ be the spectral gap of the discriminant matrix $D(P_{\beta})$ associated with $P_{\beta}$. Then, there exists a quantum algorithm that prepares the ground state of $-D$ up to $\varepsilon$ accuracy in $\ell_{2}$-norm with run time $\widetilde{\mathcal{O}}(\delta_{\textup{min}}^{-1/2}\log(1/\varepsilon))$ where $\delta_{\textup{min}}=\inf_{0\leq\beta^{\prime}\leq\beta}\delta(\beta^{\prime})$.
@@
|$\ket{\pi^{(0)}}=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}\ket{x}$
This quantum state is ground state of $D$ at $\beta=0$ since for $\beta=0$, Glauber dynamics is equivalent to hypercube walk. Similar to MIS case, given |$\ket{\pi^{(k-1)}}$, we can prepare |$\ket{\pi^{(k)}}$ up to $\varepsilon$ accuracy efficiently. The number of calls to the ground state projector by the fixed point amplitude amplification from |$\ket{\pi^{(k-1)}}$ to |$\ket{\pi^{(k)}}$ is $\widetilde{\mathcal{O}}(|\langle\pi^{(k-1)}|\pi^{(k)}\rangle|^{-1})$. The overlap can be calculated as
$\left|\!\braket{\pi^{(k-1)}}{\pi^{(k)}}\!\right|$ | $=\sum_{x\in G}\frac{\exp(-\beta_{k-1}H(x)/2)}{\sqrt{Z_{k-1}}}\frac{\exp(-\beta_{k}H(x)/2)}{\sqrt{Z_{k}}}$
@@
if we set $(\beta_{k}-\beta_{k-1})=\frac{1}{\|H\|}$. Hence, starting from |$\ket{\pi^{(0)}}$ we can prepare a schedule that maintains constant overlap with the subsequent state.

### v2

@@ Appendix B Technical Details for MaxCut Hamming and MaxBisection
$\mathbb{E}_{\pi}H(x)=\sum_{i<j}e_{ij}\mathbb{E}_{\pi}\frac{1}{2}(1-x_{i}x_{j})=\sum_{i<j}e_{ij}\Pr[x_{i}\neq x_{j}].$
Note that there are $n\choose k$ bitstrings of Hamming weight $k$. If $x_{i}\neq x_{j}$, we have the freedom to place k1k-11-1”s in n2n-2 spots. Adding a factor of two since the same can be done for “+1+1”s, we get
@@
$\frac{\mathbb{E}_{\pi}H(x)}{n}=\frac{p}{4}(1+o(1))$ (43)
@@
Fix a graph $G$, and consider an arbitrary $x\in\{-1,1\}^{n}$ such that $\lvert x\rvert=k$ corresponding to number of +1+1’s. Swaps only occur between $x_{i}$ and $x_{j}$ that are different. All one step transitions denoted by $\sim$ are implied to be under the transposition walk.
@@
For a given fixed $x\in\{-1,1\}^{n}$ satisfying $\lvert x\rvert=k$, we know there are k(nk)k(n-k) pairs of indices such that $x_{i}\neq x_{j}$, $\binom{n-k}{2}$ indices where $x_{i}=x_{j}=-1$, and $\binom{k}{2}$ indices where $x_{i}=x_{j}=1$. Thus we can decompose the expectation as follows:
@@
$e_{ij}e_{rs}\mathbb{E}_{y\underset{P}{\sim}x}\left[\frac{1-y_{i}y_{j}}{2}\frac{1-y_{r}y_{s}}{2}\right],$
@@
    $\pm 1:\mathbb{E}_{y\sim x}\left[\frac{1-y_{i}y_{j}}{2}\frac{1-y_{r}y_{s}}{2}\right]=\frac{4}{(n-k)k}$
@@
$\mathbb{E}_{y\underset{P}{\sim}x}[\hat{y}|x_{i}=x_{j}=1]=\frac{2}{k}$
$\mathbb{E}_{y\underset{P}{\sim}x}[\hat{y}|x_{i}=x_{j}=-1]=\frac{2}{n-k}.$
Then we compute $H(x)[2\mathbb{E}_{y\sim x}[H]-H(x)]$:
@@ Appendix C Constrained Short Path via Penalized Objective
The Markov chain $\mathcal{M}=(\mathcal{X},P,\pi)$ is the random walk on the hypercube, and hence $\pi$ is the uniform distribution over $\{-1,1\}^{n}$. Thus, the short path Hamiltonian is the same as [DPCB23] but with the penalized cost Hamiltonian.
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be random walk on the $n$-bit Hamming cube. Let $H:\left\{-1,1\right\}^{n}\rightarrow\mathbb{R}$ be a diagonal Hamiltonian with ground state energy
@@
If the number of constraints is $m=\Theta(n)$, then there exists a short path algorithm with runtime
@@
The penalty method ensures that our algorithm only outputs feasible solutions, but the framework only provides a speedup over unconstrained brute-force search. Hence, the generalized short path algorithm is significantly more effective.
@@ D.2 Ground State Preparation for Glauber Mixer
[…]
@@
Let $\delta(\lambda)>0$ be the spectral gap of the discriminant matrix $D(P_{\lambda})$ associated with $P_{\lambda}$. Then, there exists a quantum algorithm that prepares the ground state of D-D up to $\varepsilon$ accuracy in $\ell_{2}$ norm with run time $\widetilde{\mathcal{O}}(\delta_{\textup{min}}^{-1/2}\log(1/\varepsilon))$ where $\delta_{\textup{min}}=\inf_{0\leq\lambda^{\prime}\leq\lambda}\delta(\lambda^{\prime})$.
@@
[…]
$|\braket{\sqrt{\pi^{(0)}}|\sqrt{\pi^{(1)}}}|=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{\lambda_{1}^{1/2}}{\sqrt{Z_{1}}}\geq\frac{\sqrt{n\lambda_{1}}}{\sqrt{n\lambda_{1}+\sum\limits_{|\mathcal{I}|>1}\lambda_{1}^{\mathcal{|I|}}}}$ $=\Omega(1),$
[…]
$\left|\!\braket{\sqrt{\pi}^{(k-1)}|\pi^{(k)}}\!\right|$ $=\sum_{x\in\mathcal{X}}\frac{\lambda_{k-1}^{|x|/2}}{\sqrt{Z_{k-1}}}\frac{\lambda_{k}^{|x|/2}}{\sqrt{Z_{k}}}$
@@
[…]
@@
Let $\delta(\beta)>0$ be the spectral gap of the discriminant matrix $D(P_{\beta})$ associated with $P_{\beta}$. Then, there exists a quantum algorithm that prepares the ground state of D-D up to $\varepsilon$ accuracy in $\ell_{2}$-norm with run time $\widetilde{\mathcal{O}}(\delta_{\textup{min}}^{-1/2}\log(1/\varepsilon))$ where $\delta_{\textup{min}}=\inf_{0\leq\beta^{\prime}\leq\beta}\delta(\beta^{\prime})$.
@@
[…]
$\left|\!\braket{\sqrt{\pi^{(k-1)}}|\sqrt{\pi^{(k)}}}\!\right|$ $=\sum_{x\in G}\frac{\exp(-\beta_{k-1}H(x)/2)}{\sqrt{Z_{k-1}}}\frac{\exp(-\beta_{k}H(x)/2)}{\sqrt{Z_{k}}}$
@@
[…]